# The $C^{0,1}$ Itô-Ventzell formula for weak Dirichlet processes

### v3

@@ 5.1 Formulas for quadratic variations
Let $W$ be a one-dimensional Brownian Motion, $X=M+A$ be a continuous weak Dirichlet FQV process, and $f:\Omega\times[0,T]\times\mathbb{R}\to\mathbb{R}$ be such that $f(\cdot,x)$ is previsible for all $x\in\mathbb{R}^{d}$ fixed, and continuously differentiable in $x$ for all $t\in[0,T]$. Moreover, we assume that for all $K\subset\mathbb{R}$ compact exists a $C_{K}(\omega)\geq 0$ with $\mathbb{E}[C_{K}]<\infty$ such that $\sup_{(t,x)\in[0,T]\times K}(|f(\omega,t,x)|+|f_{x}(\omega,t,x)|)\leq C_{K}(\omega)$. Then, it holds that
@@ 5.2 A representation result for strong solutions of SPDEs
We say that $F$ is a strong solution to the stochastic elliptic problem (5.1) if there exists a sequence $(F^{n})_{n\in\mathbb{N}}$ with local characteristics $(\beta^{n}(\omega,t,x),\gamma^{n}(\omega,t,x))$ and a sequence $(h^{n})_{n\in\mathbb{N}}$ such that for every $n\in\mathbb{N}$ and a.e. $\omega\in\Omega$ holds:
@@
    $\lim_{n\to\infty}\int_{0}^{t}(F_{x}(r,X_{r})-F^{n}_{x}(r,X_{r}))\mathrm{d}^{-}A_{r}$ $+\int_{0}^{t}\mathrm{tr}((\gamma_{x}(r,X_{r})-\gamma^{n}_{x}(r,X_{r}))\sigma(r,X_{r}))\mathrm{d}r$
    $+\int_{0}^{t}(F_{x}(r,X_{r})-F^{n}_{x}(r,X_{r}))b(r,X_{r})\mathrm{d}r=0.$ (5.2)

### v4

@@ 5.1 Formulas for quadratic variations
Let $W$ be a one-dimensional Brownian Motion, $X=M+A$ be a continuous weak Dirichlet FQV process, and $f:\Omega\times[0,T]\times\mathbb{R}\to\mathbb{R}$ be such that $f(\cdot,x)$ is previsible for all $x\in\mathbb{R}^{d}$ fixed, $f(t,W_{t})$ is a.s. càdlàg, and $f$ is continuously differentiable in $x$ with the derivative being Hölder-continuous (for an $\alpha>0$) for all $t\in[0,T]$. Moreover, we assume that for all $K\subset\mathbb{R}$ compact exists a $C_{K}(\omega)\geq 0$ with $\mathbb{E}[C_{K}^{2}]<\infty$ such that $\sup_{(t,x)\in[0,T]\times K}(|f(\omega,t,x)|+|f_{x}(\omega,t,x)|)\leq C_{K}(\omega)$. Then, it holds that
@@ 5.2 A representation result for strong solutions of SPDEs
We say that a stochastic flow $F$ with local characteristics $(\beta({t},x),\gamma({t},x))$ is a strong solution to the stochastic elliptic problem (5.1) if there exists a sequence $(F^{n})_{n\in\mathbb{N}}$ with local characteristics $(\beta^{n}(\omega,t,x),\gamma^{n}(\omega,t,x))$ and a sequence $(h^{n})_{n\in\mathbb{N}}$ such that for every $n\in\mathbb{N}$ and a.e. $\omega\in\Omega$ holds:
@@
    $\lim_{n\to\infty}\Big{|}\int_{0}^{t}(F_{x}(r,X_{r})-F^{n}_{x}(r,X_{r}))\mathrm{d}^{-}A_{r}\Big{|}$ $+\Big{|}\int_{0}^{t}\mathrm{tr}((\gamma_{x}(r,X_{r})-\gamma^{n}_{x}(r,X_{r}))\sigma(r,X_{r}))\mathrm{d}r\Big{|}$
    $+\Big{|}\int_{0}^{t}(F_{x}(r,X_{r})-F^{n}_{x}(r,X_{r}))b(r,X_{r})\mathrm{d}r\Big{|}=0.$ (5.2)
    In particular, all of these integrals are assumed to exist.